87. Find the next number in the series:
1, 7, 21, 46, 85, 141,?
A. 207
B. 217
C. 227
D. 237

Answer by flame8855(424)

1, 7, 21, 46, 85, 141
series has differences
6, 14, 25, 39, 56,
which has differences
8, 11, 14, 17,
11-8 = 3
so 17 +3 =20
20+56 = 76
141+76 = 217
so the answer is B. 217